OpenStudy (anonymous):

Suppose w = [x/y] + [y/z], x = e2 t, y = 2 + sin( 1 t), z = 2 + cos( 3 t ). A. Use the chain rule to find [dw/dt] as a function of x, y, z, and t. Do not rewrite x, y, and z in terms of t, and do not rewrite e2 t as x. [dw/dt] = ??? Note: Use exp() for the exponential function. Your answer should be an expression in x, y, z, and t; e.g. "3x - 4y" B. Use part A to evaluate [dw/dt] when t = 0. ???

OpenStudy (turingtest):

x=(e^2)t or x=e^(2t) or x=2et ?

OpenStudy (anonymous):

e^(2t)

OpenStudy (turingtest):

I'm starting of with \[dw/dt=[y(dx/dt)-x(dy/dt)]/y^2+[z(dy/dt)-y(dz/dt)]/z^2\] (not the final answer) is this the kind of thing that looks like what you've been doing in class? And what do you mean by exponential as "exp0"?